# A Parallel CPU-GPU Framework for Cost-Bounded DFS with Applications to IDA* and BTS

### v1

@@ Abstract
The rapid advancement of GPU technology has unlocked powerful parallel processing capabilities, creating new opportunities to enhance classic search algorithms. A recent successful application of GPUs is in compressing large pattern database (PDB) heuristics using neural networks while preserving heuristic admissibility. However, very few algorithms have been designed to exploit GPUs during search. Several variants of A* exist that batch GPU computations. In this paper we introduce a method for batching GPU computations in depth first search. In particular, we describe a new cost-bounded depth-first search (CB-DFS) method that leverages the combined parallelism of modern CPUs and GPUs. This is used to create algorithms like Batch IDA*, an extension of the Iterative Deepening A* (IDA*) algorithm, or Batch BTS, an extensions of Budgeted Tree Search. Our approach builds on the general approach used by Asynchronous Parallel IDA* (AIDA*), while maintaining optimality guarantees. We evaluate the approach on the 3x3 Rubik’s Cube and 4x4 sliding tile puzzle (STP), showing that GPU operations can be efficiently batched in DFS. Additionally, we conduct extensive experiments to analyze the effects of hyperparameters, neural network heuristic size, and hardware resources on performance.
@@ Introduction
In recent years, there has been exponential growth in computational resources, particularly in GPUs (Dally, Keckler, and Kirk 2021; Rotem et al. 2022). Modern CPUs also continue to evolve with enhanced parallel processing capabilities, enabling faster execution of complex algorithms. Simultaneously, GPUs have become indispensable for computation-intensive tasks due to their massive parallelism, capable of performing millions of operations simultaneously. This rapid advancement in both CPU and GPU technologies has been well exploited in many fields of Artificial Intelligence, especially deep learning (Schrittwieser et al. 2020; Yao et al. 2024).
How can classic search algorithms such as Iterative Deepening A* (IDA*) (Korf 1985) and more modern algorithms like Budgeted Tree Search (BTS) (Helmert et al. 2019; Sturtevant and Helmert 2020) benefit from this growth in technology? Several approaches (Li et al. 2022; Zhou and Zeng 2015; Agostinelli et al. 2019, 2024) have been developed to enhance search algorithms using the parallel processing capabilities of modern GPUs. Li et al. (2022) introduced the first admissible neural network heuristic that compresses a large pattern database (PDB) heuristic with less information loss than standard compression techniques (Felner et al. 2007; Helmert, Sturtevant, and Felner 2017). Batch A* (Li et al. 2022) was then designed to use the admissible heuristic to find optimal solutions. Batch A* is a variant of the A* algorithm (Hart, Nilsson, and Raphael 1968), collecting states into batches to evaluate their heuristics through a single neural network lookup. This technique, called batch heuristic evaluations, utilizes GPU parallelism and significantly improves performance over performing individual neural network lookups for each state. However, Batch A* remains slower than the plain A* algorithm that uses PDB heuristics. Agostinelli et al. (2019) introduced the Batch Weighted A* (BWAS) algorithm, which combines weighted A* search (Pohl 1970; Ebendt and Drechsler 2009) with batch heuristic evaluations.
Because algorithms like BTS and IDA* are both designed around cost-bounded depth-first search (CB-DFS), the aim of this paper is to design a parallel CPU/GPU implementation of CB-DFS to efficiently use neural network heuristics during search. We demonstrate CB-DFS inside Batch IDA* has two distinct phases working together: search and heuristic evaluation. The search phase is modified from the AIDA* algorithm (Reinefeld and Schnecke 1994), utilizing CPU parallelism to asynchronously generate and explore different subtrees in CB-DFS. Meanwhile, the heuristic evaluation phase continuously collects nodes generated by all subtrees into batches, which are then processed by the heuristic model on the GPU.
We then prove that using CB-DFS inside Batch IDA* still guarantees the discovery of optimal solutions. In our experimental setting, IDA* and BTS behave identically, so we only formulate and evaluate Batch IDA*. We show that Batch IDA* outperforms Batch A* and performs comparably to AIDA* in the Rubik’s Cube domain. In addition, we show that scaling the number of GPUs accelerates Batch IDA* by enabling simultaneous evaluations.
Related Work
In recent years, several parallel versions of the IDA* algorithm have been introduced. AIDA* (Reinefeld and Schnecke 1994), a highly parallel iterative-deepening search algorithm, was designed for large-scale asynchronous Multiple Instruction, Multiple Data (MIMD) systems. The algorithm partitions the search space and processes it asynchronously across multiple CPU processors. Taking a different approach, Horie and Fukunaga (2017) investigate the parallelization of IDA* on GPUs using a block-based approach. The proposed Block-Parallel IDA* (BPIDA*) assigns subtrees to blocks of threads that execute on the same streaming multiprocessors (SMs). BPIDA* takes advantage of local shared memory of within a SM to reduce warp divergence and improve load balancing.
GPU parallelism has also been effectively utilized in other search algorithms. Zhou and Zeng (2015) introduced the first parallel variant of the A* search algorithm called GA* that leverages the computational power of GPUs. GA* uses multiple parallel priority queues to manage the Open list, enabling the simultaneous extraction and expansion of nodes across GPU threads. The heuristic computations are also parallelized across the GPU cores to optimize performance further. GA* is up to 45x faster than a traditional CPU-based A* implementations in large and complex search spaces. Q* search (Agostinelli et al. 2024) employs deep Q-networks to calculate combined transition costs and heuristic values for child nodes in a single forward pass, thereby eliminating the need to explicitly generate them.
Edelkamp and Sulewski (2009) investigate bitvector-based search algorithms, where the GPU’s parallel processing capabilities allow for the efficient handling of state expansion and duplicate detection. The GPU is also employed for ranking and unranking permutations, computing hash functions, and managing the search frontier, all of which are parallelized to exploit the GPU’s architecture. Meanwhile, other approaches have formulated A* and weighted A* algorithms as differentiable and end-to-end trainable neural network planners (Yonetani et al. 2021; Archetti, Cannici, and Matteucci 2021). These data-driven approaches, rather than learning the heuristic function, take a raw image as an input and convert it to a guidance map by an encoder. GPUs have also been instrumental in learning heuristics. For example, Li et al. (2022) developed an admissible heuristic for the sliding tile puzzle (STP) and TopSpin using an ensemble of neural networks and classifier quantiles. Other approaches (Agostinelli et al. 2021; Arfaee, Zilles, and Holte 2011; Thayer, Dionne, and Ruml 2011; Pándy et al. 2022) have also focused on heuristic learning, though without guaranteeing admissibility.
Background
@@
In heuristic search, the task is to find a path in a graph $\{G=\{V,E\},s,g,c,h\}$ from a start state $s\in V$ to a goal state $g\in V$, where $c:E\rightarrow\mathbb{R}^{+}$ is a cost function associated with the edges between states. The heuristic function $h(v)$ provides an estimate of the distance from a state $v$ to the goal $g$. The heuristic is considered admissible if, for all states $v$, $h(v)$ does not exceed the true shortest distance $h^{*}(v)$ to the goal. It is consistent if, for any two states $a$ and $b$, the heuristic satisfies $h(a)\leq c(a,b)+h(b)$. In large state spaces, the graph $G$ is represented implicitly, meaning it is generated online by expanding states and exploring their neighbors. A* and IDA* are guaranteed to find optimal solutions when the heuristic is admissible.
@@
Pattern Database (PDB) heuristics are widely utilized in problems that exhibit exponential growth. These heuristics abstract the original graph $V$ into a reduced state space $\phi(V)$. In this abstract space, edges between vertices in the original graph are preserved in the abstracted graph, meaning if an edge exists between $v_{1}$ and $v_{2}$ in $V$, a corresponding edge will exist between $\phi(v_{1})$ and $\phi(v_{2})$ in $\phi(V)$. As a result, abstract distances are admissible estimates of distances in $V$. PDBs can reduce the size of the state space exponentially with only a small loss in heuristic accuracy (Felner, Sturtevant, and Schaeffer 2009).
@@
Standard PDB compression methods (Felner et al. 2007; Helmert, Sturtevant, and Felner 2017) primarily treat the PDB as a table of numbers. These methods group entries to reduce the PDB’s size and replace each entry in a group with the smallest value in that group to ensure the heuristic remains admissible. One common method is DIV, where k adjacent entries are grouped by dividing the index by k. Another method is MOD, which combines entries offset by $\frac{m}{k}$ in a PDB with $m$ total entries using the modulo operator.
@@
Neural networks have been used to compress PDB heuristics. ADP (Samadi et al. 2008) utilizes a diverse range of techniques to uphold admissibility. This includes a unique loss function to penalize overestimation more than underestimation, utilizing a decision tree to partition states into smaller groups, and employing ANNs at the end nodes of the decision tree, which only undergo training on subsets of the state space. Any states with inadmissible heuristics are then stored in a hash table. The batch algorithms described in this paper are orthogonal to the use ADP for heuristics, but ADP would likely need to be redesigned to work efficiently on modern GPU architectures.
Li et al. (2022) propose treating the learning of heuristics as a classification problem rather than using regression. They use two other techniques to guarantee admissibility. The first is to adjust the quantile of the distribution to select heuristic values by leveraging the ordered nature of classes. For instance, a quantile of 0.5 returns the first class that has a cumulative probability of 0.5 from small to large classes. By selecting a lower quantile, the network can generate conservative estimates that are less likely to overestimate the cost to the goal. The second is to use an ensemble of neural networks. Each network in the ensemble generates its own heuristic estimate, and the minimum of these estimates is used as the final heuristic value. This technique leverages the diversity of the ensemble to produce an admissible heuristic.
A*, IDA* and BTS
@@
IDA* combines a cost-bounded depth-first search (CB-DFS) with iterative deepening to find optimal solutions. IDA* performs a series of depth-first searches, each with an increasing cost threshold, which is determined by the current path cost and the heuristic estimate to the goal. The algorithm starts with an initial threshold set to the heuristic estimate of the start node. It explores all paths in a depth-first order with costs less than or equal to this threshold. If the goal is not found, the threshold is increased to the minimum cost that exceeds the current threshold. This process is repeated until the goal is reached. IDA* is memory-efficient, as it only requires storage for the current path. However, if the number of nodes in each iteration does not grow fast enough, the cost of the iterations may outweigh the cost of the final iteration, increasing the total expansions from $O(N)$ to $O(N^{2})$. BTS Helmert et al. (2019) is a variant of IBEX which is identical to IDA* when the iterations grow by a suitable constant factor, but reduces the worst-case overhead to $O(N\log C^{*})$, where $C^{*}$ is the optimal solution cost. As in IDA*, BTS repeatedly invokes CB-DFS, just with more aggressive thresholds and node expansion limits. Thus, both IDA* and BTS can be improved through GPU parallelization of CB-DFS.
AIDA*
AIDA* parallelizes the CB-DFS portion of IDA* in three key phases (as shown in Figure 1):
  • Initial Data Partitioning: In this phase, all processors work in parallel to redundantly expand the initial few levels of the search tree using an iterative-deepening approach. This generates sufficient nodes to ensure that each processor has enough work for the next phase. At the end of this phase, duplicate nodes are eliminated from the frontier nodes.
  • Distributed Node Expansion: Each processor is assigned a set of nodes from the first phase, expanding these nodes further to create a larger set. At the end of this phase, the local node arrays of the individual processors make up the global work queue used for dynamic load balancing in the third phase. Each work packet in the work queue is the root of a subtree to be explored in the next phase.
@@
  • Asynchronous Search: In this phase, processors independently and asynchronously explore different subtrees of the search space using an iterative-deepening strategy. Each processor continues working on its assigned subtree until either a solution is found or the subtree is fully explored. Since the work queue is shared among all threads, idle threads dynamically retrieve unprocessed work packets from the queue as they complete their current assignments, ensuring continuous utilization of all processing resources.
@@
GPUs use a hierarchical execution model that enhances parallel processing efficiency. At the heart of this model is the kernel, which executes across multiple thread blocks. Each block contains warps, groups of threads that execute the same instructions in parallel. Threads within a block can share data through on-chip memory, but blocks themselves work independently without direct communication.
@@
CPUs excel at tasks with complex control flows and high instruction-level parallelism, making them ideal for sequential tasks. In contrast, GPUs are optimized for handling large, independent data sets, which makes them especially effective for neural network computations. In convolutional neural networks (CNNs), for example, a weight matrix and a subset of inputs are assigned to a thread block, where threads compute convolution outputs by sharing data through fast on-chip memory. Since different convolution operations are independent of each other, they can be executed in parallel batches across multiple thread blocks (Jeon 2023). This data-level parallelism gives GPUs a significant advantage over CPUs in neural network training and inference.
Batch A*
Batch A* (Li et al. 2022) is a modification of A* that allows it to delay node heuristic evaluation when using neural network heuristics. When nodes are generated, they are placed into a batch queue instead of the open list. Once the batch queue is full, the heuristics of the states in the batch are evaluated in parallel, and then they are inserted into the A* open list for search to continue.
Problem Definition
@@
In a heuristic search problem, the heuristic, $h$, can come from any source. In this paper we are interested in the neural heuristic search problem. In this problem, $h$ is a neural heuristic. This means that $h\in H_{NN}$, where $H_{NN}$ is the set of all heuristics that are computed by a neural network. In recent work, heuristics in $H_{NN}$ have been learned from PDB heuristics (Li et al. 2022), but we do not distinguish the source of the heuristic or how it was trained. Experimental results evaluate the quality of heuristics we currently have access to, but we are working under the assumption that neural network heuristics will continue to improve, and thus we should build and evaluate algorithms that can be efficiently used with such heuristics during search.
Method
We now introduce the SingleGPU Batch IDA* algorithm, provide its pseudocode, and prove its theoretical guarantees. Then, we discuss the MultiGPU Batch IDA* algorithm. These algorithms are based on a CPU/GPU implementation of CB-DFS. Batch BTS has a different outer loop, but can use CB-DFS with an additional expansion limit. Since Batch IDA* and Batch BTS perform identically in our setting, we present the ideas using the simpler Batch IDA* setting. But, our approach parallelizes the common CB-DFS used by both, not the top-level control of IDA* or BTS.
@@
Algorithm 2 Generate works for parallel CB-DFS
1:function GenerateWork($s$,$d_{init}$,history)
2:     if size(history)=$d_{init}$ then
3:         work $\leftarrow$ Work($s$)
4:         work.init $\leftarrow$ history
5:         works.add(work)
6:         return
7:     end if
8:     actions $\leftarrow$ env.GetActions($s$)
9:     for each action in actions do
10:         history.add(action)
11:         $s$leftarrow env.ApplyAction($s$,action)
12:         GenerateWork($s$,$d_{init}$,history)
13:         $s$leftarrow env.UndoAction($s$,action)
14:         history.pop()
15:     end for
16:end function
@@
9:     while miss $<$ workNum do             $\triangleright$ wait for all works to
@@
SingleGPU Batch IDA*
SingleGPU Batch IDA* extends AIDA* by leveraging batch heuristic evaluations using a learned heuristic $h_{M}$. The algorithm builds the CB-DFS tree in a manner consistent with AIDA*. It begins by replicating the first two phases of AIDA* to generate a Work Queue for parallel processing. Each piece of work in the Work Queue represents a subtree rooted at a node generated during the initial phase. The remaining procedure of the algorithm is generally depicted in Figure 2. In each CB-DFS call, every thread retrieves several pieces of work from the Work Queue and stores them in a stack. Threads process each piece of work in depth-first order, generating new nodes as they traverse its subtree. Once a piece of work is completed—meaning all nodes in its subtree have been generated—the thread takes on a new one. CB-DFS completes when all threads have processed the work in their stacks under the current cost threshold and the Work Queue is empty.
The newly generated nodes from all threads are collected in a vector called batch. When the batch reaches its maximum size, a batch-processing thread manages it in three steps: copying the batch from the CPU to the GPU, performing heuristic lookup using the model $h_{M}$, and copying the heuristic evaluations back to the CPU. After assigning the evaluations to their nodes, the batch is cleared, allowing the next set of generated nodes to be processed.
The detailed steps of SingleGPU Batch IDA* are outlined in Algorithm 1. The first two phases are executed via the GenerateWork function (line 5). This function is presented in Algorithm 2, which constructs the CB-DFS tree up to a depth of $d_{init}$. The third phase is managed by invoking the CB-DFS function in parallel across available CPU threads at each cost threshold (line 8). The cost threshold is updated (UpdateThreshold function) after the CB-DFS with the current threshold is completed (line 10).
The CB-DFS function (provided in Algorithm 3) has one primary change compared to CB-DFS function in AIDA*. In AIDA*, each thread processes one piece of work at a time and does not begin another until the current one is fully completed. However, applying this method to Batch IDA* limits the maximum batch size:
$batch\ size\leq\max_{s\in S}b(s)*n,$ (1)
where $b(s)$ is the branching factor of state $s$ and $n$ is the number of CPU threads. This limitation occurs because heuristic evaluations for successors aren’t immediately ready, forcing the CB-DFS to wait for the completion of batch evaluation. This is restrictive, as large batch sizes are needed to fully utilize GPU parallelism. To overcome this, each thread needs to process multiple pieces of work altogether (Algorithm 3). For this purpose, each thread maintains a work stack, with the number of pieces specified by workNum (lines 6-8). This approach allows for larger batches by adjusting workNum while maintaining the same number of threads.
@@
The DoIteration function (Algorithm 4) expands subtrees one at a time. While executing the DoIteration function, if a subtree cannot progress further—meaning the heuristic evaluation is not yet available for the top node (line 5)— the thread switches to the next subtree in its stack. Upon revisiting the same subtree, the thread resumes from the same node if its heuristic evaluation is ready; otherwise, it switches again to the next subtree.
The batch processing is done via the ProcessBatch function, as shown in Algorithm 5. In this function, we consider a timeout which invokes batch processing even if the batch is not full. This mechanism helps when only a few pieces of work remain and the number of generated nodes is insufficient to fill the batch.
@@
8:         if $h_{M}$($s$,goal)$<$bound then
@@
14:         $s_{next}$leftarrow env.ApplyAction($s$, action)
@@
Algorithm 5 Process of heuristic batch
1:function ProcessBatch()
2:     wait for the batch to get full or for the timeout to expire
3:     copy batch from CPU to GPU
4:     results $\leftarrow$ $h_{M}$(batch)
5:     copy results from GPU to CPU
6:     return results
7:end function

### v2

# A Parallel CPU-GPU Framework for Batching Heuristic Operations in Depth-First Heuristic Search

@@ Abstract
The rapid advancement of GPU technology has unlocked powerful parallel processing capabilities, creating new opportunities to enhance classic search algorithms. This hardware has been exploited in best-first search algorithms with neural network-based heuristics by creating batched versions of A* and Weighted A* that delay heuristic evaluation until sufficiently many states can be evaluated in parallel on the GPU. But, research has not addressed how depth-first algorithms like IDA* or Budgeted Tree Search (BTS) can have their heuristic computations batched. This is more complicated in a tree search, because progress in the search tree is blocked until heuristic evaluations are complete. In this paper we show that GPU parallelization of heuristics can be effectively performed when the tree search is parallelized on the CPU while heuristic evaluations are parallelized on the GPU. We develop a parallelized cost-bounded depth-first search (CB-DFS) framework that can be applied to both IDA* and BTS, significantly improving their performance. We demonstrate the strength of the approach on the 3x3 Rubik’s Cube and the 4x4 sliding tile puzzle (STP) with both classifier-based and regression-based heuristics.
@@ Introduction
There has been significant recent growth in computational resources, particularly in GPUs (dally2021evolution; rotem2022intel). GPUs have become indispensable for computation-intensive tasks due to their massive parallelism, capable of performing millions of operations simultaneously. Modern CPUs also continue to evolve with enhanced parallel processing capabilities, enabling faster execution of complex algorithms. This advancement in both CPU and GPU technologies has been well exploited in many fields of Artificial Intelligence, especially deep learning (schrittwieser2020mastering; yao2024tree).
In classical search algorithms, several approaches (li2022optimal; zhou2015massively; agostinelli2019solving; agostinelli2024q) have been developed to enhance search using the parallel processing capabilities of modern GPUs. One potential use of deep learning on GPUs is to learn heuristics to guide search. For instance, li2022optimal introduced admissible neural network heuristics that compresses a large pattern database (PDB) heuristic with less information loss than standard compression techniques (felner2007compressed; helmert2017variable).
A* (hart1968formal) and Weighted A* (pohl1970heuristic) have had batch versions developed which can use GPU-based heuristics more efficiently for optimal (li2022optimal) and sub-optimal search (agostinelli2019solving). The batched versions collect states into batches to evaluate their heuristics in a single parallel neural network lookup using the GPU. This technique, called batch heuristic evaluations, utilizes GPU parallelism and significantly improves performance over performing individual neural network lookups for each state.
Algorithms like IDA* (korf1985depth) have not yet had batched variants built. The depth-first nature of IDA* search complicates the batching process, because in a depth-first search there are often only a few states available at one time, while batching is most efficient when hundreds of states are available for computing heuristic values in parallel. The same issue applies to algorithms such as Budgeted Tree Search (BTS) (DBLP:conf/ijcai/HelmertLLOS19; sturtevant2020btsguide), as both IDA* and BTS are built upon performing repeated cost-bounded depth-first searches (CB-DFS).
This paper addresses the issue by observing that approaches used for CPU parallelization of CB-DFS make many more states available for batching, and thus enable efficient use of GPU-based heuristics. By re-designing CB-DFS for parallel search, we can then effectively build batch versions of IDA* and BTS. The effectiveness of the batched versions of IDA* and BTS is shown using both regression-based and classifier-based heuristics on the Rubik’s Cube and 15-puzzle domains. Batching is highly effective in improving the performance of these algorithms, resulting in over a $40\times$ improvement in search speed. This improvement in search performance opens the door for further research on improving heuristic quality, and for designing new search algorithms to handle the inadmissible values that we expect to find in large neural network-based heuristics.
Background and Related Work
@@
In heuristic search, the broad task is to find a path in a graph $\{G=\{V,E\},s,g,c,h\}$ from a start state $s\in V$ to a goal state $g\in V$, where $c:E\rightarrow\mathbb{R}^{+}$ is a cost function associated with the edges between states. The heuristic function $h(v)$ provides an estimate of the distance from a state $v$ to the goal $g$. The heuristic is considered admissible if, for all states $v$, $h(v)$ does not exceed the true shortest distance $h^{*}(v)$ to the goal. It is consistent if, for any two states $a$ and $b$, the heuristic satisfies $h(a)\leq c(a,b)+h(b)$. In large state spaces, the graph $G$ is represented implicitly, meaning it is generated online by expanding states and exploring their neighbors. A* and IDA* are guaranteed to find optimal solutions when the heuristic is admissible (felner2011inconsistent).
@@
Pattern Database (PDB) heuristics (culberson1998pattern) and the related merge and shrink framework (helmert2007flexible) are widely utilized, particularly in problems that exhibit exponential growth (Gnad_Sievers_Torralba_2023). These heuristics abstract the original graph $V$ into a reduced state space $\phi(V)$. In this abstract space, edges between vertices in the original graph are preserved in the abstracted graph, meaning if an edge exists between $v_{1}$ and $v_{2}$ in $V$, a corresponding edge will exist between $\phi(v_{1})$ and $\phi(v_{2})$ in $\phi(V)$. As a result, abstract distances are admissible estimates of distances in $V$. PDBs can reduce the size of the state space exponentially with only a small loss in heuristic accuracy (felner2009abstraction).
@@
Standard PDB compression methods (felner2007compressed; helmert2017variable) primarily treat the PDB as a table of numbers. These methods group entries to reduce the PDB’s size and replace each entry in a group with the smallest value in that group to ensure the heuristic remains admissible. One common method is DIV, where k adjacent entries are grouped by dividing the index by $k$. Another method is MOD, which combines entries offset by $\frac{m}{k}$ in a PDB with $m$ total entries using the modulo operator.
@@
Neural networks have been used to compress PDB heuristics. ADP (samadi2008compressing) used a range of techniques to ensure admissibility. These included a unique loss function to penalize overestimation, a decision tree to partition states, and employing ANNs only for the resulting subsets of states. Any states with inadmissible heuristics were then stored in a hash table. ADP was developed prior to current hardware and is orthogonal to work in this paper on designing more efficient algorithms that use these heuristics.
li2022optimal also studied approaches for learning admissible heuristics. They treated the learning of heuristics as a classification problem rather than using regression, because in NP-complete problems the solution length is polynomial, meaning a small number of heuristic values (classes), while the state space is exponential. Admissibility is guaranteed in two ways. First, because the heuristic classes are ordered, the classification quantile used for the predicted class can be adjusted to ensure admissibility. Second, an ensemble of neural networks can be trained, with the minimum value from the ensemble used as the prediction. This technique leverages the diversity of the ensemble to produce an admissible heuristic.
Search Algorithms
A variety of search algorithms can be used for solving shortest path problems.
@@
IDA* combines a cost-bounded depth-first search (CB-DFS) with iterative deepening to find optimal solutions. IDA* performs a series of depth-first searches, each with an increasing cost threshold, which is determined by the current path cost and the heuristic estimate to the goal. IDA* increases the cost threshold conservatively so it can terminate once the goal is found. IDA* is memory-efficient, as it only requires storage for the current path. However, if the number of nodes in each iteration does not grow fast enough, the sum of costs of the iterations may outweigh the cost of the final iteration, increasing the total expansions to $O(N^{2})$.
BTS (DBLP:conf/ijcai/HelmertLLOS19) is identical to IDA* when the iterations grow by a suitable constant factor, but reduces the worst-case overhead to $O(N\log C^{*})$, where $C^{*}$ is the optimal solution cost. As in IDA*, BTS repeatedly invokes CB-DFS, just with more aggressive thresholds and additional node expansion limits. Thus, both IDA* and BTS can be improved through GPU parallelization of CB-DFS.
AIDA* is designed to parallelize the CB-DFS portion of IDA*. While some forms of depth-first search cannot be parallelized well (reif1985depth), AIDA* exploits the structure of the problem to create independent subproblems. AIDA* has three phases for CPU parallelization of the CB-DFS portion of IDA*, as shown in Figure 1:
Initial Data Partitioning
In this phase a single search is performed from the root of the search tree where the leaves of the search have cost greater than the current cost threshold. The threshold is increased until there are sufficiently many leaves above the cost threshold to efficiently perform the following algorithmic steps. At the end of this phase, duplicate nodes are eliminated from the frontier nodes.
Distributed Node Assignment
Next, the leaf nodes from the data partitioning are put into a shared work queue for distributed assignment of work. Path information is maintained to prevent the search from returning to the parents.
Distributed CPU Search
@@
In this phase, processors independently search the subtrees found below the states in the shared work queue, maintaining the lowest unexpanded $f$-cost in the search. In each iteration either a solution is found and the search terminates, or no solution is found, and the search is repeated at the next cost threshold. Since the work queue is shared among all threads, idle threads dynamically retrieve unprocessed work from the queue as they complete their current work.
@@
GPU parallelization is fundamentally different than CPU parallelization. While in CPU parallelization we need independent subtasks, GPU parallelization works best with correlated tasks that are solved with the same instructions but different data. GPUs use a hierarchical execution model that enhances parallel processing efficiency. At the heart of this model is the kernel, which executes across multiple thread blocks. Each block contains warps, groups of threads that execute the same instructions in parallel. Threads within a block can share data through on-chip memory, but blocks themselves work independently.
@@
GPU Parallelization of Algorithms
In recent years, several parallel versions of the IDA* algorithm have been introduced. AIDA* (reinefeld1994aida), a highly parallel iterative-deepening search algorithm, was designed for large-scale asynchronous Multiple Instruction, Multiple Data (MIMD) systems. The algorithm partitions the search space and processes it asynchronously across multiple CPU processors. Taking a different approach, horie2017block investigate the parallelization of IDA* on GPUs using a block-based approach. The proposed Block-Parallel IDA* (BPIDA*) assigns subtrees to blocks of threads that execute on the same streaming multiprocessors (SMs). BPIDA* takes advantage of local shared memory of within a SM to reduce warp divergence and improve load balancing.
GPU parallelism has also been effectively utilized in other search algorithms. zhou2015massively introduced the first parallel variant of the A* search algorithm called GA* that leverages the computational power of GPUs. GA* uses multiple parallel priority queues to manage the Open list, enabling the simultaneous extraction and expansion of nodes across GPU threads. The heuristic computations are also parallelized across the GPU cores to optimize performance further. GA* is up to 45x faster than a traditional CPU-based A* implementations in large and complex search spaces. Q* search (agostinelli2024q) employs deep Q-networks to calculate combined transition costs and heuristic values for child nodes in a single forward pass, thereby eliminating the need to explicitly generate them.
edelkamp2009parallel investigate bitvector-based search algorithms, where the GPU’s parallel processing capabilities allow for the efficient handling of state expansion and duplicate detection. The GPU is also employed for ranking and unranking permutations, computing hash functions, and managing the search frontier, all of which are parallelized to exploit the GPU’s architecture. Meanwhile, other approaches have formulated A* and Weighted A* algorithms as differentiable and end-to-end trainable neural network planners (yonetani2021path; archetti2021neural). These data-driven approaches, rather than learning the heuristic function, take a raw image as an input and convert it to a guidance map by an encoder. GPUs have also been instrumental in learning heuristics. For example, li2022optimal developed an admissible heuristic for the sliding tile puzzle (STP) and TopSpin using an ensemble of neural networks and classifier quantiles. Other approaches (agostinelli2021obtaining; arfaee2011learning; thayer2011learning; pandy2022learning) have also focused on heuristic learning, though without guaranteeing admissibility.
CB-DFS for Neural Heuristics
@@
In a heuristic search problem, the heuristic, $h$, can come from any source. This paper studies the neural heuristic search problem. In this problem, $h$ is a neural heuristic. This means that $h\in H_{NN}$, where $H_{NN}$ is the set of all heuristics that are computed by a neural network. In recent work, heuristics in $H_{NN}$ have been learned from PDB heuristics (li2022optimal), and general techniques for heuristic learning have been described (khandelwal2024towards). The aim of this paper is to improve algorithms that make use of such heuristics. Experimental results evaluate the quality of heuristics we currently have access to, but we are working under the assumption that neural network heuristics will continue to improve, and thus improved algorithms will be broadly beneficial.
We now introduce the SingleGPU Batch CB-DFS algorithm, which can be used with both IDA* to create Batch IDA* and BTS to create Batch BTS. We provide pseudocode and prove theoretical properties. Then, we discuss the MultiGPU CB-DFS algorithm.
@@
9:  while miss << workNum do $\triangleright$ wait for all works to
@@
SingleGPU Batch CB-DFS
Batch heuristic lookups in best-first search algorithms like A* are relatively easy because once the search gets started, there are many states waiting in the open list. But, in a cost-bounded depth-first tree, the depth-first search cannot continue until the $f$-cost of all children is known. Thus, to get sufficient states for batching heuristic lookups, we must either have a large enough branching factor to permit efficient batching, search speculatively beyond the cost limit, or run several CB-DFS searches in parallel so we can batch states across subtrees. This paper explores the last approach.
The overall structure of SingleGPU CB-DFS is illustrated in Figure 2. It begins similarly to AIDA* by generating a Shared Work Queue with subtrees that can be searched independently. However, instead of having each CPU take a single subtree, the CPUs now fill a work stack with multiple subtrees. A single expansion is performed on the first subtree, which results in a set of children that need their heuristics evaluated before the search can continue. These states are placed into a shared heuristic batch queue on the CPU, and then work continues on the next subtree in the work stack, generating more states until the heuristic batch queue is full. If multiple CPUs are available, they can execute this process in parallel.
Once the heuristic batch queue is full, a dedicated batch-processing thread manages the batch evaluation of states. This involves (1) copying the data to GPU memory, (2) evaluating the batch on the GPU using the model $h_{M}$, and then (3) returning the resulting heuristics to each of the CPUs. The batch is then cleared, allowing the next set of generated states to be processed. More than one batch is maintained in RAM so the CPUs can continue to search in parallel to the batch evaluation. The CB-DFS completes when the shared work queue is exhausted and all threads have completed any searches remaining in their work stacks. This completes the high-level CB-DFS.
The pseudo-code showing batch CB-DFS and how it is integrated into SingleGPU Batch IDA* is outlined in Algorithm 1. The GenerateWork function (line 5) is called once before the main search loop begins, instructing an initial search tree to depth $d_{init}$. BatchIDA* then invokes CB-DFS in parallel across CPU threads given the cost threshold of the high-level search (line 8). The cost threshold is updated (UpdateThreshold function) after the CB-DFS with the current threshold is completed (line 10).
@@
The DoIteration function (Algorithm 3) expands subtrees one at a time. While executing the DoIteration function, if a subtree cannot progress further—meaning the heuristic evaluation is not yet available for the top node (line 5) – the thread switches to the next subtree in its stack. Upon revisiting the same subtree, the thread resumes from the same node if its heuristic evaluation is ready; otherwise, it switches again to the next subtree.
The batch processing is done via the ProcessBatch function, which is not shown. The most important detail of this algorithm is that it has a timeout after which a batch is evalauted even if it is not completely full. This is important when very little work is available at the end of an iteration, and impacts the integration with BTS.
@@
8:   if $h_{M}$($s$,goal)<<bound then
@@
14:   $s_{next}$ $\leftarrow$ env.ApplyAction($s$, action)